OpenStudy (anonymous):

On which of the following intervals is the function f(x) = 4 cos(2x - π) decreasing?

OpenStudy (anonymous):

@rishavraj do you know this?

OpenStudy (anonymous):

@tejasvir

OpenStudy (anonymous):

@rishavraj

OpenStudy (aum):

Find f'(x) Find the interval where f'(x) is negative. That will be the interval where f(x) is decreasing.

OpenStudy (anonymous):

@aum is the answer C?

OpenStudy (anonymous):

can you also help on this For an angle Θ with the point (15, -8) on its terminating side, what is the value of cosine?

OpenStudy (aum):

How did you arrive at C for the first one? Show the work and I can tell you if it is correct or not.

OpenStudy (anonymous):

someone just messaged me and told me the answer was C

OpenStudy (aum):

Sorry, I can't give out answers. Try the method suggested in my first reply.

OpenStudy (anonymous):

okay can you help me with this second question then? @aum

OpenStudy (anonymous):

For an angle Θ with the point (15, -8) on its terminating side, what is the value of cosine? I thought it'd be 15, because that's cosine but im not sure

OpenStudy (aum):



OpenStudy (anonymous):

okay so it'd be 15 / hypotenuse

OpenStudy (aum):

Since it is in the fourth quadrant, cosine will be positive. Find the length of OP. cos(t) = adjacent /hypotenuse = 15 / OP = ?

OpenStudy (anonymous):

15 / 17?

OpenStudy (aum):

yes. 15/17.03 = 0.8808

OpenStudy (anonymous):

wait, but the answer choices don't have that

OpenStudy (aum):

What are the choices?

OpenStudy (anonymous):

the choices are -8/15 15 / 18 -8 15

OpenStudy (anonymous):

any ideas??

OpenStudy (aum):

I think the choices are wrong. The second choice should have been 15/17.

OpenStudy (anonymous):

are you sure? maybe the answer is just 15?

OpenStudy (aum):

cosine of any angle can never be greater than +1 or less than -1.

OpenStudy (anonymous):

but how can the choices be wrong? could it be we're doing it wrong?

OpenStudy (aum):

Can you copy and paste the exact question?

OpenStudy (anonymous):

Wait it was a -8, maybe that's why

OpenStudy (aum):

The last link does not refer to the question we are discussing.

OpenStudy (anonymous):

I'm sorry, wrong one, give me one second

OpenStudy (aum):

You can quickly rule out the last two choices because \(\large -1 \le \cos(\theta) \le 1\). You can rule out the first choice because the point (15, -8) lies in the fourth quadrant where cosine is positive. That leaves with just choice 2 except it should be 15/17 and not 15/18. (sometimes there are typos in books and notes).

OpenStudy (aum):

Just choose choice B and most probably it will accept the answer as correct even though there is a typo.

OpenStudy (anonymous):

OKay thank you

OpenStudy (anonymous):

@aum ONE LAST ONE I PROMISE

OpenStudy (anonymous):

A heated piece of metal cools according to the function c(x) = (.5)x - 7, where x is measured in hours. A device is added that aids in cooling according to the function h(x) = -x - 2. What will be the temperature of the metal after 2 hours?

OpenStudy (anonymous):

Choices: -4° Celsius 28° Celsius 32° Celsius 38° Celsius

OpenStudy (anonymous):

c(x) = (.5)^(x-7)

OpenStudy (aum):

For the first question you can look at this graph and figure out in which interval f(x) = 4 cos(2x - π) is decreasing: http://www.wolframalpha.com/input/?i=plot++4cos%282x-pi%29%2C+ [0%2C+2pi]

OpenStudy (anonymous):

I figured that out, can you help me on this new one? @aum please I'm in a huge rush now

OpenStudy (aum):

Find c(2) and h(2) and add them up.

OpenStudy (anonymous):

I tried but when you plug in 2 for c(x) it comes out to be like 2?

OpenStudy (anonymous):

.25**

OpenStudy (aum):

What is the initial temperature of the heated metal?

OpenStudy (anonymous):

Basically I just need help finding c(2)

OpenStudy (anonymous):

It doesn't say.... lol

OpenStudy (aum):

Is c(x) = (.5)x - 7 or (.5)^x - 7 (Is x an exponent?)

OpenStudy (anonymous):

(x - 7) is the full exponent

OpenStudy (aum):

Then use parenthesis to group items and ^ to denote exponent. c(x) = (0.5)^(x-7) x = 2 c(2) = 0.5^(-5) = 32 h(x) = -x - 2 h(2) = -2 -2 = -4 32 + (-4) = 28

OpenStudy (anonymous):

so 28 is the answer??

OpenStudy (aum):

If it is not clear, (0.5)^(-5) = (1/2)^(-5) = (2/1)^(+5) = 32. Yes, 28 degrees Celsius is the answer.

OpenStudy (anonymous):

Thank you so much. Looking over my test, i actually have one more question. You don't have to answer but I'd appreciate if you checked my work

OpenStudy (anonymous):

A coach wants to know if football players who warm up before a game will score more points than those who don't warm up. He will break the team up into two groups and make sure each player gets equal time on the field. He will then count how many points each player makes during the game. What statistic should he study?

OpenStudy (anonymous):

The mean number of points earned by each player The mean number of points earned by each group The standard deviation of the number of points earned by each player The standard deviation of the number of points earned by each group

OpenStudy (anonymous):

I put D

OpenStudy (anonymous):

@aum

OpenStudy (aum):

I would go with B but I am not sure. For the first question, the function f(x) = 4 cos(2x - π) is decreasing in the interval x = pi/2 and x = pi.

OpenStudy (anonymous):

are you 100% sure? someone messaged me and said it was decreasing from pi to 3 pi /2? @aum

OpenStudy (aum):

yes.